OpenStudy (anonymous):

Given f(x)=3x^2-2x, find f(a+1)

OpenStudy (asnaseer):

just replace each 'x' in the expression with 'a+1'

mathslover (mathslover):

allow me to give an example : f(x) = x^2 + 1 then f(2) = (2^2)+1=4+1=5

mathslover (mathslover):

do it similarly .. as @asnaseer said : put x = (a+1) and then solve

OpenStudy (anonymous):

i did but idk how to square it... like do i do 3*a*^2 or what?

OpenStudy (asnaseer):

ok, so lets concentrate on just that part

mathslover (mathslover):

\[\large{3(a+1)^2-2(a+1)}\] can you tell what is (a+1)^2 ?

OpenStudy (asnaseer):

\[(a+1)^2=?\]can you do this?

OpenStudy (anonymous):

 this is the question know